# HyperGaussians: High-Dimensional Gaussian Splatting for High-Fidelity Animatable Face Avatars

### v2

@@ A.1.1 Formulation
where we used $\det\boldsymbol{\Sigma}_{a|b}=\det\boldsymbol{\Lambda}_{aa}^{-1}=\bigl{(}\det\boldsymbol{\Lambda}_{aa}\bigr{)}^{-1}$ in step $(1)$, $\det\boldsymbol{\Lambda}_{aa}=\det\boldsymbol{L}_{aa}\boldsymbol{L}_{aa}^{\top}=\bigl{(}\det\boldsymbol{L}_{aa}\bigr{)}^{2}$ in step $(2)$, and $\det\boldsymbol{L}_{aa}=\prod_{i=1}^{m}(\boldsymbol{L}_{aa})_{i,i}$ in step $(3)$. The $\log$ at step $(3)$ is applied element-wise.
Again, we can compute this quantity efficiently using the inverse covariance trick (Sec. A.1.1). These values are summed up across the conditional distributions for position, rotation, and scale. In order to render these uncertainties, we further apply a sigmoid function and map the values to colors. This agreement between the uncertainty estimates and what would intuitively be considered difficult regions emerges without explicit supervision.
A.2 Supplementary Experiments
A.2.1 Qualitative Results
We show supplementary video results for self- and cross-reenactment on a supplementary HTML page.
@@ A.2.1 Qualitative Results
In addition, Fig. 9 provides more examples for comparing the convergence speed of FlashAvatar without (top) vs. with HyperGaussians (bottom).
A.2.2 Ablation Study
We complement the ablation study from the main paper with supplementary results for different MLP configurations in Tab. 4 and Fig. 12. The default FlashAvatar MLP [51] has 6 layers with 256 neurons, totaling 375K parameters. Replacing vanilla 3D Gaussians with HyperGaussians adds optimizable parameters (see Sec. A.1.1). One might assume that simply increasing the parameter count for the FlashAvatar MLP would improve the results, but this is not the case. We ablate different MLP configurations in Tab. 4. Adding more parameters to the MLP does not perform as well as adding HyperGaussian. In fact, it performs the same while slowing down the rendering speed. FlashAvatar with vanilla 3DGS runs at 347 FPS. With a large MLP, this number drops to 158 (for $256\times 40$, 2.6 M parameters) and 178 (for $512\times 11$, 2.7 M parameters). With HyperGaussian ($n=8$ and 2.6 M parameters), the original MLP ($256\times 6$) outperforms the other MLP variants for all metrics while maintaining a rendering speed of 300 FPS. All metrics and rendering times were computed on a single Nvidia GeForce RTX 2080 Ti for images with resolution $512\times 512$. In summary, the HyperGaussians’ performance improvement cannot be matched by increasing the complexity of the MLP. HyperGaussians boost the performance while maintaining fast rendering speed.
Figure 10: Comparison with GaussianHeadAvatar on multiview videos from the NeRSemble [19] dataset. HyperGaussians (Ours) demonstrate more accurate reconstructions for complex deformations, thin structures, and specular highlights. We outperform in all metrics: higher PSNR, SSIM ($10^{-1}$), and lower LPIPS ($10^{-1}$).
A.2.3 GaussianHeadAvatar Integration
We demonstrate the broad applicability of HyperGaussians by successfully integrating them into GaussianHeadAvatar [53]. Fig. 10 shows significant visual improvements: more accurate modeling of complex skin deformations (rows 1, 2), better alignment of fine structures in hair strands (row 2), and sharper specular reflections in the eyes (row 1). These quality enhancements come with minimal computational overhead, increasing training time by only 5.6% (from 9h to 9.5h). These results confirm that HyperGaussians can effectively enhance existing Gaussian-based avatars with negligible performance impact.
Figure 11: Qualitative comparison with NDGS integrated into FlashAvatar. The limited degrees of freedom of NDGS cause misalignments of thin structures and edges. The numbers show PSNR, SSIM ($10^{-1}$), and LPIPS ($10^{-2}$).
A.2.4 Comparison with NDGS
Our approach fundamentally differs from NDGS [10] in its mathematical formulation and capabilities. While NDGS uses the conditional covariance matrix, which is independent of the latent code, to directly represent the size and shape, our HyperGaussians apply multivariate Gaussians on each attribute where the conditional means dynamically adapt the location, scale, and orientation of the derived 3D Gaussians in response to the latent code. This crucial difference gives HyperGaussians the necessary degrees of freedom to model thin geometry and complex deformations with higher accuracy. Fig. 11 shows our method eliminates the artifacts visible in NDGS (particularly on glass frames) and produces substantially more precise geometry at boundaries.
MLP (width×$\times$×depth) HyperGaussian? Latent # Param. PSNR ↑ SSIM $\bigl{(}10^{-1}\bigr{)}$ LPIPS$\bigl{(}10^{-2}\bigr{)}$
256$\times$6 no - 375 K 29.43 9.466 5.107
256$\times$40 no - 2.6 M 28.10 9.380 5.720
512$\times$11 no - 2.7 M 29.50 9.472 5.122
256$\times$6 yes 1D 1.2 M 29.73 9.492 5.066
256$\times$6 yes 2D 1.4 M 29.92 9.503 5.000
256$\times$6 yes 4D 1.8 M 29.89 9.507 4.994
256$\times$6 yes 8D 2.6 M 29.99 9.510 4.978
256$\times$6 yes 16D 4.1 M 29.92 9.511 4.978
256$\times$6 yes 32D 7.2 M 29.89 9.511 4.976
256$\times$6 yes 64D 13.4 M 29.91 9.512 5.020
Table 4: Supplementary Ablations. Simply increasing the parameter count for the FlashAvatar MLP does not improve the metrics. Our HyperGaussians, however, improve the performance of the original MLP out-of-the-box. As an additional benefit, HyperGaussians render 300 FPS while the deeper MLPs drop to 158 FPS ($256\times 40$) and 178 FPS ($512\times 11$), respectively. FlashAvatar with vanilla 3DGS renders fastest at 347 FPS. Green denotes the best and Yellow the second best.
A.3 Societal Impact
It is important to be aware that photorealistic, high-quality face avatars from monocular videos can have societal implications. While our novel HyperGaussian representation contributes to exciting possibilities for entertainment, communication, and virtual experiences, it could potentially be misused to spread misinformation and deception. Realistic face avatars could be exploited to produce convincing deepfakes, potentially undermining trust in visual media and influencing societies and politics. We strongly condemn any form of abuse or malicious use of our research and advocate for responsible development and application of face avatar technology, always in strict accordance with local laws and regulations.

### v3

@@ D.2 Comparison with NDGS
GaussianHeadAvatar (GHA) [71] 24.10 8.819 20.273
NDGS (GHA) [13] 24.22 8.820 20.235
Ours (GHA) 24.38 8.819 19.768
Table 3: Quantitative comparison with GHA and NDGS in the multi-view setting. Notice that NDGS is unable to match the improvements of HyperGaussians in terms of PSNR and LPIPS, and performs only marginally better on SSIM. This highlights the limited capabilities of NDGS due to its reduced degrees of freedom.
D.3 Ablation Study
We complement the ablation study from the main paper with supplementary results for different MLP configurations in Tab. 4 and Fig. 11. The default FlashAvatar MLP [70] has 6 layers with 256 neurons, totaling 375K parameters. Replacing vanilla 3D Gaussians with HyperGaussians adds optimizable parameters (see Sec. B.1). One might assume that simply increasing the parameter count for the FlashAvatar MLP would improve the results. However, this is not the case. We ablate different MLP configurations in Tab. 4. Adding more parameters to the MLP does not perform as well as adding HyperGaussian. In fact, it performs the same, or worse than the baseline, while significantly slowing down rendering speed. FlashAvatar with vanilla 3DGS runs at $347$ FPS. With a large MLP, this number drops to $158$ (for $256\times 40$, 2.6M parameters) and $178$ (for $512\times 11$, 2.7M parameters). With HyperGaussians ($n=8$ and 2.6M parameters), the original MLP ($256\times 6$) outperforms the other MLP variants for all metrics while maintaining a rendering speed of $300$ FPS. All metrics and rendering times were computed on a single NVIDIA GeForce RTX 2080 Ti for images with resolution $512\times 512$. In summary, the HyperGaussians’ performance improvement cannot be matched by increasing the complexity of the MLP. HyperGaussians boost the performance while maintaining fast rendering speed. Moreover, while FlashAvatar with vanilla 3DGS achieves higher FPS, we observe faster convergence with HyperGaussians. Fig. 12 visualizes the training progress over time and shows that HyperGaussians achieve a higher quality for a given time budget.
[…]
Table 4: MLP and Latent Dimensionality Ablations. Simply increasing the parameter count for the FlashAvatar MLP does not improve the metrics. Our HyperGaussians, however, improve the performance of the original MLP out-of-the-box. As an additional benefit, HyperGaussians render at around $300$ FPS while the deeper MLPs, with comparable parameter counts, drop to $158$ FPS ($256\times 40$) and $178$ FPS ($512\times 11$), respectively. Note that the drop of $10$ FPS for 1D is likely due to memory bottlenecks caused by poor cache and vector load/store locality. Green denotes the best and Yellow the second best.
We ablate the effect of different latent dimensionalities in Tab. 4. We find that HyperGaussian are robust towards different latent dimensions. A latent dimension of $8$ performs very well, but we already observe an improvement for a single latent dimension ($n=1$) over the vanilla 3DGS variant. The top row corresponds to the FlashAvatar baseline [70], which does not use any HyperGaussians.
Appendix E Societal Impact
It is important to be aware that photorealistic, high-quality face avatars from monocular videos can have societal implications. While our novel HyperGaussian representation offers exciting possibilities for entertainment, communication, and virtual experiences, it could potentially be misused to spread misinformation and deception. Realistic face avatars could be exploited to produce convincing deepfakes, potentially undermining trust in visual media and influencing societies and politics. We strongly condemn any form of abuse or malicious use of our research and advocate for responsible development and application of face avatar technology, always in strict accordance with local laws and regulations.
Figure 10: Qualitative Comparison with GHA and NDGS on subjects from the NeRSemble [28] dataset. While NDGS helps improve skin deformations, the resulting wrinkles are less pronounced (rows 2 and 4 show blurrier, washed-out results). Moreover, NDGS lacks visual fidelity for specular highlights (row 3) and struggles with accurate geometric alignment (teeth artifacts in row 1 and downward displacement of the glass frames in row 4).